[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Aug 26, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Aug 26 20:56:41 UTC 2017


HSU Shift Log for Sat Aug 26 2017 16:56:35 GMT-0400 (EDT)

Shift: H

Forecaster: L Avila

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: Caught a HU Watch in CRP's area while creating the  
21z National TCV. Danielle notified them and it was removed. Make sure you  
QC the hazard grids!
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170826/c5a6d41a/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list